# Breaking the News: A LLM-based Game where Players Act as Influencer or Debunker for Raising Awareness About Misinformation

### v1

@@ 3.1. Overview of Game Design
The game we designed focused on the challenges of managing information in a health crisis. One player assumes the role of an influencer (Player 1) hired by a company to promote a remedy based on traditional medicine philosophies but lacking extensive scientific support. This player can create and disseminate misinformation about the remedy. The second player takes on the role of a journalist-debunker (Player 2) advocating for a newly developed medicine supported by current scientific research. This player’s goal is to debunk the misinformation spread by the influencer. The game features a system that simulates public opinion, whereby an LLM models the reactions of five characters who read players’ messages. Previous studies showed the robustness of the LLM in evaluating the quality of an argument (Mirzakhmedova et al., 2024) (See details in Section 3.3). The objective is to sway the simulated public opinion in favor of one’s position.
Figure 2. Game Flow. (A) Both players read the current round’s news and instructions. (B) Player 1 starts first as an influencer to generate misinformation by choosing whether to buy the hints. (C) Player 1 inputs text and sends a request to the GPT-4o API. (D) GPT-4o API returns a response of public opinion. (E) Player 2 then starts to counter Player 1’s misinformation by choosing whether to buy the hints. (F) Player 2 inputs text and sends a request to the GPT-4o API. (G) GPT-4o API returns an updated public opinion.
@@ 3.2.1. Game Flow
Participants were randomly assignedto one of two roles: influencer (Player 1) or journalist-debunker (Player 2) or Player 2. Both players are provided with instructions, including the setting, the fundamental reality regarding the effectiveness of the two medicines, their roles and tasks, demographic information about the public personas of the people with whom they would interact, and the rules of the game (Figure 2A).
The game unfolds over four rounds, each featuring a new set of updated news. In each round, the influencer (Player 1) begins by reviewing the news and any instructions. They can also decide to buy hints provided in the game using in-game currency (Figure 2B). Once ready, the influencer types their information and it is published (Figure 2C). The LLM reacts to this information by simulating public opinion, with this impact of the information on public opinion visible to both players (Figure 2D).
Next, the journalist-debunker (Player 2) takes their turn. Player 2 reviews the current public opinion and where appropriate counters any misinformation by typing their debunking response, which they then publish (Figure 2F). If needed, they can purchase customized hints (Figure 2E) or read the instructions. After publishing their response and receiving feedback from the LLM, the round ends, and the game progresses to the next round (Figure 2G).
At the end of each round, both players can view the results, which reflect the journalist-debunker’s impact on misinformation. The process for the remaining three rounds is the same as in the first round.
@@ 3.2.2. Narratives
The game is set in a fictional small country with called Southland. Historically, Southland has been known for producing renowned medical doctors and pharmacists, and its people take pride in their traditional medical practices. However, there are ongoing debates in this country about the comparative merits of modern healthcare methods and traditional medicine. The sudden outbreak of the ”Zinc Virus” further amplifies these debates. As the healthcare system becomes overwhelmed and the scientific community unable to provide an effective treatment because of limited knowledge about this novel virus, residents turn to traditional medicine in search of hope.
@@
We set the game in a health crisis scenario because, in real life, situations marked by scientific uncertainty – where authorities can be unable to provide confident full explanations or advice – often fuels rumors and speculation about treatments(Wang et al., 2019). These dynamics were observed during the Ebola(Fung et al., 2016), Zika(Wood, 2018), and COVID-19 pandemics(Suarez-Lledo and Alvarez-Galvez, 2021). In such scenarios, traditional medicine frequently promoted to prevent or treat viruses(Lam et al., 2021; Mutombo et al., 2023). Additionally, we incorporated the traditional medicine controversy into our narrative as these debates are well-known to our participants, who were of an East Asian background. Research showed that 48.4% of Hong Kong residents reported using traditional medicine before the COVID-19 pandemic(Lam et al., 2021). Similarly, a national survey in South Korea found a 74.8% prevalence of traditional medicine use overall(Ock et al., 2009). In China, traditional medicine is formally integrated into the healthcare system as a widely practised modality(Chung et al., 2023). This cultural relevance provided a familiar backdrop for the game’s content, allowing players to engage intuitively and to craft responses based on their lived experiences.
@@
To avoid biases or stereotypes associated with traditional medicine, we use ”Product R” to refer to traditional medicine. Before gameplay, we explained to players the nature of Product R, particularly stating that it has not undergone rigorous testing and lacks scientific consensus on its safety. We also introduce another medicine called ”Max” to represent science-based medicine, and we also explain the nature of Max before the game begins. Players are informed that Max has been subject to rigorous clinical trials, with the results published in a peer-reviewed medical journal, however, these studies have demonstrated inconsistent effectiveness against the Zinc virus. We deliberately avoided making vaccines a topic of this game to prevent players’ pre-existing attitudes toward vaccines from influencing their behaviour in the game.
@@
To create the ”News” pieces in the game, we first researched examples of misinformation by reviewing relevant literature. We selected key features of misinformation and incorporated them into the game’s events (See details in 2.1). To ensure the misinformation was portrayed realistically, we investigated real-world examples from fact-checking websites [^1] https://www.snopes.com/, reputable news outlets [^2] https://www.wsj.com/, and medical websites [^3] https://www.webmd.com/ [^4] https://sciencebasedmedicine.org/ [^5] https://healthfeedback.org/. For instance, we represented ”less credible sources” using personal stories, viral videos, and newspapers that can objectively be classified as being biased. In representing biased newspapers, we are further inspired by the political orientation that may influence preferences for medical care. For example, research has shown that voters who tend to support anti-corruption parties are less likely to seek services from mainstream healthcare providers, and are more inclined to use alternative treatments(Valtonen et al., 2023). Similarly, the ”creating fake news for profit” feature was illustrated through scenarios where investors leveraged expert endorsements to boost sales. Next, we crafted ”News” pieces for the game based on these findings. However, any similarities to real-world sites or companies are purely coincidental; all names and events was invented solely for the purpose of this study. Table 1 summarizes the key features of misinformation identified in the literature, the associated cited papers, and how these are reflected in the in-game news. The full version of the ”News” is in the Appendix. Lastly, To ensure balanced gameplay, we used a ChatGPT-4o model to review the narrative and provide an opinion on the difficulty to players of dealing with each piece of news in the game context. Taking this opinion into account, we made further revisions and corrections when we conducted two pilot tests with four people. The goal was to ensure the game was balanced and gave opportunities to win the game for both players.
Table 1. Key misinformation features and corresponding ”News” in the Game
@@
False information is often shared by lower-quality media. However, political ideology is positively associated with individuals’ perceptions of media trustworthiness. As a result, people may place greater trust in biased sources that align with an individual’s political partisanship.(Zhang and Ghorbani, 2020; Molina et al., 2021; Hanley et al., 2023) A newspaper reported that a renowned medical expert has advocated Product R, claiming its herbal ingredients could potentially treat the Zinc Virus. This newspaper is known for its anti-corruption stance. (Round 1)
Personal, negative, and opinionated tones predominate in misinformation narratives which frequently provoke dread, anxiety, and mistrust of institutions(Bessi et al., 2015; Porat et al., 2019). A widow shared her husband’s experience. She suspects that Max was ineffective and believes it may have caused renal impairment, eventually leading to her husband’s death. She claims, “He was given a medication we demanded he NOT receive, and his health quickly went downhill,” ultimately resulting in him being “on a ventilator working most of the time at 100%.” (Round 2)
Fake news, some of which is purposely fabricated to cause harm, generate financial returns, or spread confusion(Mustafaraj and Metaxas, 2017; Hanley et al., 2023; Shu et al., 2017) A journalist discovered that the institution of traditional medicine where the famous medical expert works received significant funding from billionaire Jack. Additionally, Jack’s ex-wife owns a company that produces and promotes traditional medicine products like R. (Round 3)
False rumors will create feedback loops and evolve into more intense and extreme versions over time(Shin et al., 2018) A popular short video claims that a doctor who practiced alternative medicine and R was murdered to protect the profits of “Big Pharma”. More people are attracted to believe in the validity of traditional medicines and advocate for their use while opposing new drugs. Growth in sentiment that resistance to traditional medicine amounts to being an attack on their cultural heritage. (Round 4) This round’s advocate for R becomes more intense than in Round 1, with the focus being less on effectiveness but rather patriotic sentiment.
@@ 3.2.3. Instructions and Hints
To support players, the game includes instructional content that features definitions(Molina et al., 2021; Wu et al., 2019), examples(Goodman and Carmichael, 2020), and strategies for both creating and debunking misinformation. This content draws on insight gained into misinformation from research literature and practice. For the misinformation creator role, we applied the Elaboration Likelihood Model and used simple examples to teach players to craft persuasive misinformation(Petty et al., 1984; Moran et al., 2016). For the misinformation debunker role, we used an Agence France-Presse fact-checking style-guide and an guide published by the EU on communicating with proponents of conspiracy theories (Agence France-Presse, 2023; European External Action Service, 2021). These were the inspiration for a user-friendly guide we developed to assist game players to identify misinformation and equip them with effective debunking strategies. Additionally, the game offers two hints during each round to inspire players. The detailed hint is crafted by the authors using the same materials used to create the in-game news. The simple hint is generated by ChatGPT-4o model. When authors used it to review the news and ensure balanced gameplay, it provided concise suggestions on how each player might respond from their perspective. (Full instructions and hints can be found in the supplementary materials)
@@ 3.2.4. LLM Basis
We implemented the LLMs to play the role of ”public opinion” in the game for three reasons. Firstly, LLMs perform well when processing dynamic natural language(Hu et al., 2024). Secondly, LLMs demonstrate memory capability, such as with working memory being applied to the context of a conversation, and long-term memory allowing past conversational information to be taken into account(Hu et al., 2024). Research also demonstrates that generative agents powered by LLMs build a high degree of capability for responding to the context of a conversation(Park et al., 2023b; Yin et al., 2024). Thirdly, LLMs excel at role-playing tasks in the game. Research indicated(Wang et al., 2023)(Hu et al., 2024) that directly inserting natural language descriptions of a role’s identity enable LLMs to make better quality evaluations in conversational tasks. These capabilities allow LLMs to effectively serve as ”evaluators” in the game, generating continuous, context-aware dialogue and feedback. This approach proves more effective than traditional prebuilt game mechanics, such as trigger keywords for assessment.
3.2.5. Game Mechanics
The player’s success depends on the reaction of LLM-simulated public opinion. Inspired by Jaffe et al.’s research on game balancing systems(Jaffe et al., 2012), we use a metric to determine which player’s information is more effective in term of it being more compelling and convincing to the public.
In each round, each LLM character evaluates the messages provided by the players and provides an opinion and trust level score on a 10-point scale:
  • 10: The character fully trusts the misinformation published by the Influencer (Player 1)
  • 0: The character completely trusts the debunking response published by the Journalist-debunker (Player 2)
  • If the final score is above 5, Influencer (Player 1) wins.
  • If the final score is below 5, Journalist-debunker (Player 2) wins.
The game includes a reward system based on the trust score scale, and this serves to maintain player engagement. Players earn in-game currency based on the average public trust score, and can use this to purchase game hints.
@@ 3.2.5. Game Mechanics
3.2.6. Game Interface
The game’s interface is text-based, thus making it easy for players to generate information, with the assistance of references and instructions. The interface is divided into sections:
  • Information Viewing Section (Figure 3A): This section includes news and all previously published information from both players. It is designed to allow players to quickly review existing information and adjust their in-game strategies accordingly.
  • Public Opinion Section (Figure 3B): This section displays the LLM-simulated public opinion, represented visually by five circles and an area labeled Average Public Opinion Score, which shows the overall level of trust for the information. The data is retrieved from the OpenAI API. Each circle is color-codedby trust score, ranging from blue (0), indicating complete distrust, to red (10), indicating complete trust in Player 1’s information. The average public opinion score is calculated from the trust scores of the five characters. Players can click on each circle to view detailed reactions from individual characters.
  • Text Editing Section (Figure 3C): This section provides tools for text editing, including instructions and purchasable hints that help players in generating their information. These hints vary each round, both in content and cost, and are designed to help players generate more convincing information.
  • In-Game Currency Panel (Figure 3D): This panel displays the current player’s available in-game currency, helping them to decide whether or not to buy hints.
@@ 3.3. Prompt Engineering
We employed the GPT-4o model to be the LLM-simulated role in the game. The prompt in the game is structured into five sections (Figure 4A): Story Context, Role Descriptions and Rules (Criteria), Personas, Response Format and Additional Notes. These five sections form the System message (Figure 4B).
3.3.1. Story Context
This section (Figure 4A-1) serves as the basis of the game. It provides an LLM generated model of the fundamental setting for the ”world”. This includes background information and each round’s news. The model has to evaluate every player’s actions based on this predefined ”world” setting. The full example prompt appears below:
”The Southland, with a 6,000-year history, boasts rich natural resources, diverse ecosystems, and a culture that values liberty, free expression, and media independence. Historically, Southland has been known for its renowned doctors and pharmacists, and its people are proud of their traditional medicine …”
3.3.2. Role Description
The ”Role Descriptions and Rules (Criteria)” section (Figure 4A-2) serves to ensure the model understands its role in the game, i.e., how the game runs. A example prompt is below:
Your role is to simulate how the five personas react to each piece of information they receive …
Additionally, when evaluating and generating public opinions, we provide specific rules and criteria for the model to follow (Figure 4A-2.2, 2.3, 2.4, 2.5, 2.6). Specifically, we explain to the model of how to evaluate each player’s inputs separately for Player 1 (Figure 4A-2.2) and Player 2 (Figure 4A-2.3), as well as how to evaluate each persona’s reactions to the misinformation (Figure 4A-2.4) and anti-misinformation (Figure 4A-2.5) input by each player. Lastly, we provide criteria by which the model provides trust level scores (Figure 4A-2.6).
3.3.3. Personas
@@
We designed five distinct fictional personas, each with unique demographic characteristics (Figure 4A-3). These personas were developed based on insights from the literature, incorporating four key group factors: demographics, psychological traits, personality, and behavioral features. These factors guided how each persona responded to misinformation and anti-misinformation messages (Nan et al., 2022; Liu et al., 2023; Lee et al., 2024; Shin et al., 2024). (See detailed prompt in Supplementary materials)
Alex Smith. Age: 36. Gender: Male. Project Manager. Education Level: Undergraduate. Political party affiliation: Strongly support Liberal. Enjoys social activities but also values alone time for work and personal projects. Generally trusting and cooperative, but can be skeptical of new information. Balances responsibilities but can be hasty in decision-making. Generally calm and composed, rarely anxious or stressed. Open to new experiences but sometimes prefers convenience over exploration. Busy with work responsibilities and managing projects, often skims through news during short breaks. Follows news via quick-read apps and social media. Often shares articles based on headlines without reading fully. Some critical thinking skills but lacks depth in media literacy.
3.3.4. Response Format
This section defines the standard response data format to ensure the game can process the data easily after they are received (Figure 4A-4). When the data are returned to the game, a predefined framework extracts them and assigns them to the relevant game elements for display. This standardized formatting ensures the integration of the API responses into the game, allowing for consistent and accurate gameplay experiences.
Persona 1 Basic Details (e.g., age, gender, occupation), Persona Reaction, Trust Level Score, Persona 2 […], Persona 3 […], Persona 4 […], Persona 5 […]
Every time the API receives new information from players, it will return a response in this format, consisting of five personas’ public opinions. Each persona’s public opinion is composed of the demographic information, a trust level score, and reaction. In the game, these three aspects of data are extracted and displayed (Figure 5).
3.3.5. Additional Notes
This section is designed for supplementary notes such as edge cases (Figure 4A-5).
Figure 5. Extraction of OpenAI Response Data and Displayed in the Game.

### v2

@@ 3.1. Overview of Game Design
The game we designed focused on the challenges of managing information in a health crisis. One player assumes the role of an Influencer hired by a company to promote a remedy based on traditional medicine but lacking extensive scientific support. This player can create and disseminate misinformation about the remedy. The second player takes on the role of a Journalist advocating for a newly developed medicine supported by scientific research. The goal of this player is to debunk/disprove the misinformation spread by the Influencer. The game features a system that simulates public opinion, whereby an LLM models the reactions of five characters who read players’ messages. The objective is to sway the simulated public opinion in favor of one’s position.
@@ 3.2.1. Game Flow
Participants were randomly assigned to one of two roles: Influencer or Journalist. Both players are provided with instructions, including the setting, the fundamental reality regarding the effectiveness of the two medicines, their roles and tasks, demographic information about the characters who represent public opinion, and the rules of the game (Figure 2A).
The game unfolds over four rounds, each featuring a new set of updated news. In each round, the Influencer begins by reviewing the news and any instructions. They can also decide to buy hints provided in the game using in-game currency (Figure 2B). Once ready, the Influencer types their information and it is published (Figure 2C). The LLM reacts to this information by simulating public opinion, with this impact of the information on public opinion visible to both players (Figure 2D).
Next, the Journalist takes their turn. Journalist reviews the current public opinion and where appropriate counters any misinformation by typing their debunking response, which they then publish (Figure 2F). If needed, they can purchase customized hints (Figure 2E) or read the instructions. After publishing their response and receiving feedback from the LLM, the round ends, and the game progresses to the next round (Figure 2G).
At the end of each round, both players can view the results, which reflect the Journalist’s impact on misinformation. The process for the remaining three rounds is the same as in the first round.
Figure 2. Game Flow. (A) Both players read the current round’s news and instructions. (B) Influencer starts first to generate misinformation by choosing whether to buy the hints. (C) Influencer inputs text and sends a request to the GPT-4o API. (D) GPT-4o API returns a response of public opinion. (E) Journalist then starts to counter Influencer’s misinformation by choosing whether to buy the hints. (F) Journalist inputs text and sends a request to the GPT-4o API. (G) GPT-4o API returns an updated public opinion.
@@ 3.2.2. Narratives
The game is set in a fictional small country with called Southland. Historically, Southland has been known for producing renowned medical doctors and pharmacists. However, there are ongoing debates in this country about the comparative merits of modern healthcare methods and traditional medicine. The sudden outbreak of the “Zinc Virus” further amplifies these debates. As the healthcare system becomes overwhelmed and the scientific community unable to provide an effective treatment because of limited knowledge about this novel virus, residents turn to traditional medicine in search of hope.
@@
We set the game in a health crisis scenario because, in real life, situations marked by scientific uncertainty, where authorities can be unable to provide confident full explanations or advice – often fuels rumors and speculation about treatments(Wang et al., 2019). These dynamics were observed during the Ebola(Fung et al., 2016), Zika(Wood, 2018), and COVID-19 pandemics(Suarez-Lledo and Alvarez-Galvez, 2021). In such scenarios, traditional medicine frequently promoted to prevent or treat viruses(Lam et al., 2021; Mutombo et al., 2023). Additionally, we incorporated the traditional medicine controversy into our narrative as these debates are well-known to our participants, who were of an East Asian background. Research showed that 48.4% of Hong Kong residents reported using traditional medicine before the COVID-19 pandemic(Lam et al., 2021). Similarly, a national survey in South Korea found a 74.8% prevalence of traditional medicine use overall(Ock et al., 2009). In China, traditional medicine is formally integrated into the healthcare system as a widely practised modality(Chung et al., 2023).
@@
We use “Product R” to represent traditional medicine and “Max” for science-based medicine. Before gameplay, we explained to players the nature of Product R, particularly stating that it has not undergone rigorous testing and lacks scientific consensus on its safety. Players are also informed that Max has been subject to rigorous clinical trials, with the results published in a peer-reviewed medical journal, however, these studies have demonstrated inconsistent effectiveness against the Zinc virus. We deliberately avoided making vaccines a topic of this game to prevent players’ pre-existing attitudes toward vaccines from influencing their behaviour in the game.
@@
To create the “News” pieces in the game, we first researched examples of misinformation by reviewing relevant literature. We selected key features of misinformation and incorporated them into the game’s events (See details in 2.1). To ensure the misinformation was portrayed realistically, we investigated real-world examples from fact-checking websites [^1] https://www.snopes.com/, reputable news outlets [^2] https://www.wsj.com/, and medical websites [^3] https://www.webmd.com/ [^4] https://sciencebasedmedicine.org/ [^5] https://healthfeedback.org/. For instance, we represented “less credible sources” using personal stories, viral videos, and newspapers that can objectively be classified as being biased. The portrayal of biased newspapers draws on research showing that political orientation can influence medical preferences. For example, research has shown that voters who tend to support anti-corruption parties are less likely to seek services from mainstream healthcare providers, and are more inclined to use alternative treatments(Valtonen et al., 2023). Next, we crafted “News” pieces for the game based on these findings. However, any similarities to real-world sites or companies are purely coincidental; all names and events was invented solely for the purpose of this study. Table 1 summarizes the key features of misinformation identified in the literature, the associated cited papers, and how these are reflected in the in-game news. The full version of the “News” is in the supplementary materials. Lastly, To ensure balanced gameplay, we used a GPT-4o model to review the narrative and provide an opinion on the difficulty to players of dealing with each piece of news in the game context. Taking this opinion into account, we made further revisions and corrections when we conducted two pilot tests with four people. The goal was to ensure the game was balanced and gave opportunities to win the game for both players.
Table 1. Key misinformation features and corresponding “News” in the Game
@@
False information is often shared by lower-quality media. Political ideology, however, shapes people’s perceptions of trustworthiness.(Zhang and Ghorbani, 2020; Molina et al., 2021; Hanley et al., 2023) A newspaper reported that a renowned medical expert has advocated Product R, claiming its herbal ingredients could potentially treat the Zinc Virus. This newspaper is known for its anti-corruption stance. (Round 1)
Personal, negative, and opinionated tones predominate in misinformation narratives which frequently provoke dread, anxiety, and mistrust of institutions.(Bessi et al., 2015; Porat et al., 2019) A widow shared her husband’s experience. She suspects that Max was ineffective and believes it may have caused renal impairment, eventually leading to her husband’s death. She claims, “He was given a medication we demanded he NOT receive, and his health quickly went downhill,” ultimately resulting in him being “on a ventilator working most of the time at 100%.” (Round 2)
Fake news, some of which is purposely fabricated to cause harm, generate financial returns, or spread confusion.(Mustafaraj and Metaxas, 2017; Hanley et al., 2023; Shu et al., 2017) A Journalist discovered that the institution of traditional medicine where the famous medical expert works received significant funding from billionaire Jack. Additionally, Jack’s ex-wife owns a company that produces and promotes traditional medicine products like R. (Round 3)
False rumors will create feedback loops and evolve into more intense and extreme versions over time.(Shin et al., 2018) A popular short video claims that a doctor who practiced alternative medicine and R was murdered to protect the profits of “Big Pharma”. More people are attracted to believe in the validity of traditional medicines and advocate for their use while opposing new drugs. Growth in sentiment that resistance to traditional medicine amounts to being an attack on their cultural heritage. (Round 4) This round’s advocate for R becomes more intense than in Round 1, with the focus being less on effectiveness but rather patriotic sentiment.
@@ 3.2.3. Instructions and Hints
To support players, the game includes instructional content that features definitions(Molina et al., 2021; Wu et al., 2019), examples(Goodman and Carmichael, 2020), and strategies for both creating and debunking misinformation. This content draws on insight gained into misinformation from research literature and practice. For the Influencer, we applied the Elaboration Likelihood Model and used simple examples to teach players to craft persuasive misinformation(Petty et al., 1984; Moran et al., 2016). For the Journalist, we used an Agence France-Presse fact-checking style-guide and an guide published by the EU on communicating with proponents of conspiracy theories (Agence France-Presse, 2023; European External Action Service, 2021). These were the inspiration for a user-friendly guide we developed to assist game players to identify misinformation and equip them with effective debunking strategies. Additionally, each round offers two hints to support players. The detailed hint is crafted by the authors using the same materials as the “News”. The simple hint is generated by GPT-4o model. When authors used it to review the news and ensure balanced gameplay, it provided concise suggestions on how each player might respond from their perspective. (Full instructions and hints can be found in the supplementary materials)
@@ 3.2.4. LLM Basis
We implemented the LLMs to play the role of “public opinion” in the game for three reasons. Firstly, LLMs perform well when processing dynamic natural language(Hu et al., 2024). Secondly, LLMs demonstrate memory capability, such as with working memory being applied to the context of a conversation, and long-term memory allowing past conversational information to be taken into account(Hu et al., 2024). Research also demonstrates that generative agents powered by LLMs build a high degree of capability for responding to the context of a conversation(Park et al., 2023; Yin et al., 2024). Thirdly, LLMs excel at role-playing tasks in the game. Research indicated(Wang et al., 2023)(Hu et al., 2024) that directly inserting natural language descriptions of a role’s identity enable LLMs to make better quality evaluations in conversational tasks. These capabilities allow LLMs to effectively serve as “evaluators” in the game, generating continuous, context-aware dialogue and feedback. This approach proves more effective than traditional prebuilt game mechanics, such as trigger keywords for assessment.
3.2.5. Game Mechanics and Interface
The text-based game interface having four sections that allows players to view the information clearly for decision-making during gameplay. Players can view the current round’s news and all previously published information in the Information Viewing Section (Figure 3A); they can also see the LLM-simulated public opinion in the Public Opinion Section (Figure 3B). In the Text Editing Section (Figure 3C), players can edit their information, view instructions, and purchase hints. Player success is determined by the reaction of LLM-simulated public opinion, measured through trust level scores. This scoring system helps players adjust their strategies in subsequent rounds. To maintain engagement, the game features a reward system tied to trust scores: players earn in-game currency each round based on their average score, which can be used to purchase hints. Full details of the scoring and currency system are provided in the Appendix.
@@ 3.3. Prompt Engineering
We employed the GPT-4o model and integrated prompt design techniques informed by prior work. The prompt was structured into four sections:
  • First, we provided a game story context (Figure 4-A-1) to establish a consistent narrative foundation. This ensured that all outputs generated by the LLM remained coherent and aligned with the game’s world setting.
  • Second, we assigned the LLM a specific role (Figure 4-A-2), clearly defining its responsibilities and tasks. As the core section of the prompt, we included detailed instructions for the LLM. To guide the model effectively, we applied the Rails approach(Amatriain, 2024) by predefining rules to constrain the LLM’s output. Additionally, we used the Chains approach(Amatriain, 2024) to structure the workflow clearly to let the LLM process the task step-by-step.
@@
  • Third, since the core game mechanic involves the LLM generating diverse public opinions, we used the Expert prompting approach(Amatriain, 2024) and designed five distinct fictional personas (Figure 4-A-3). This helped the LLM create content that matches each persona’s perspective. We further enriched the personas by incorporating insights from the literature and focused on four key group factors: demographics, psychological traits, personality, and behavioral features. These factors guided how each persona responded to misinformation and anti-misinformation messages.(Nan et al., 2022; Liu et al., 2023; Lee et al., 2024; Shin et al., 2024).
  • In section 4, we reinforced output consistency and quality by including explicit output format rules and example prompts (Figure 4-A-4). These examples taught the LLM(Amatriain, 2024) how to produce coherent, high-quality outputs aligned with the game design.
A more detailed version of the prompt structure can be found in the appendix .